# Modified mean field ansatz for charged polarons in a Bose-Einstein condensate

### v1

@@ Abstract
Ionic Bose polarons are quantum entities emerging from the interaction between an ion and a Bose-Einstein condensate (BEC), featuring long-ranged interactions that can compete with the gas healing length. This can result in strong interparticle correlations and enhancement of gas density around the ion. One possible approach to describe this complex system with high accuracy relies on numerical treatment such as the quantum Monte Carlo (QMC) techniques. Nevertheless, it is computationally very expensive and does not easily allow to study the system dynamics. On the other hand, a mean-field based variational ansatz in the co-moving frame can capture a sizeable change in the gas density. We apply it to the case of regularized ion-atom potential and find that it qualitatively reproduces the full numerical results. In addition, we also study the system of two pinned ions, focusing on their effective interaction induced by the bath. This approach seems to be promising for studying transport and nonequilibrium dynamics of charged (bi)polarons in condensed media.
@@
Figure 1: The bare (solid lines) interaction given by eq. (10) and the effective potential (dashed lines) resulting from eq. (6) for several atom-ion scattering length values with $c=0.0023R^{\star}$, $n_{0}^{-1/3}=R^{\star}$ and $a_{\mathrm{B}}=0.0269R^{\star}$. The cartoon illustrates how the buildup of the density of repulsive bosons leads to the dressing of the potential.
@@
$\begin{split}H_{\mathrm{LLP}}&=\frac{1}{4m_{\mathrm{I}}}\left(\bm{p}_{0}-\sum_{i}\hat{\bm{p}}_{\mathrm{B},i}\right)^{2}+\frac{\hat{P}^{2}}{m_{\mathrm{I}}}+\sum_{i,j}V_{\mathrm{IB}}(\hat{\bm{x}}_{\mathrm{B},ij})\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\sum_{i}\frac{\hat{\bm{p}}_{\mathrm{B},i}^{2}}{2m_{\mathrm{B}}}+\frac{1}{2}\sum_{i\neq j}V_{\mathrm{BB}}(\hat{\bm{x}}_{\mathrm{B},i}-\hat{\bm{x}}_{\mathrm{B},j}),\end{split}$ (7)
@@
Figure 2: Left: energy of a single polaron in units of $E^{\star}=1/[m_{\mathrm{B}}(R^{\star})^{2}]$ as a function of the regularization parameter $b$, with $n_{0}=0.1458\,(R^{\star})^{-3}$. Results of the modified mean field theory in the co-moving frame (straight red line) are compared with QMC calculation (green dots) as well as standard mean-field theory where $E_{1}=-\pi^{2}n/b$ (light blue dashed) and the GPE calculation (yellow dots). Right: results for higher density $n_{0}=(R^{\star})^{-3}$ with the T-matrix results of Christensen et al. (2021) (green dashed) and the mean-field prediction (light blue dashed) for comparison.
@@
The minimization of the MMF functional (4) yields the ground state energy of the system which we show in Fig. 2 and compare to the ladder approximation result of Christensen et al. (2021). The two approaches agree well in the limit of very weak interactions $b\gtrsim 1$ as expected. For deeper potentials, the MMF curve bends towards lower energies due to the buildup of atomic density around the ion which increases the integrated interaction energy. The standard mean field formula $E_{1}=4\pi na/m$ Scazza et al. (2022) with $a$ being the ion-atom scattering length lies very closely to the ladder approximation result as it also assumes a homogeneous gas. We also include a comparison with a QMC calculation for 192 bosons at a lower density $n_{0}=0.1458\,(R^{\star})^{-3}$. The MMF and QMC agree very well in the whole range of $b$ considered, while standard mean field theory leads to a higher estimation of energy. Also the solution of standard GPE equation for two coupled fields representing the ion and the atoms, shown as yellow dots, can only reproduce the QMC and MMF results for large values of $b$ where essentially all approaches agree with each other.
Refer to captionRefer to caption
Figure 3: Upper: the gas density as a function of the distance from the impurity $r$ for several values of $b$. The dots indicate the results of solving the functional in full three-dimensional space. Lower: the number of bosons in the polaronic cloud as a function of the potential regularization parameter $b$ for one ion (light blue) and two pinned ions at two different distances $R$ (straight and dashed red line). The other parameters are the same as in Fig. 1.
@@
Refer to captionRefer to captionRefer to caption
Figure 4: Left: Cut of the gas density profile along $z$ for $r=0$ for two pinned ions. Here $\xi_{2}=R^{\star}$, $c=0.0023R^{\star}$, $b=0.572R^{\star}$, and $n_{0}=1.9531(R^{\star})^{-3}$. The red dots indicate the results of a full three-dimensional calculation. Middle: absolute value of the system energy in units of $E^{\star}$ for two ions as a function of their separation for the same parameters and several $b$ values, right: the resulting induced effective interaction potentials. Dotted line indicates the power law derived within the ladder approximation  Ding et al. (2022) described by eq. (12).
In the case of two impurities we set up a 2D rectangular mesh, composed of 100 elements in each dimension, and second-order Lagrange polynomials for space discretization as well. Note that if we assumed $\xi_{1}=R^{\star}$, the corresponding value for the healing length is now $\xi_{2}=(\sqrt{3}/2)R^{\star}$ for the same condensate density $n_{0}$. The impurity-impurity-boson correlation density profiles, depicted in Fig. 4, reach a maximum value where the impurities are located. We further observe that the number of bound bosons is roughly twice the single-impurity case for both potentials, as presented by the red curve in the lower panel of Fig. 3. This would be expected if the induced interactions are relatively weak such that the system can be regarded as two almost independent polarons. Indeed, inspecting the density profile shows that the gas density is quite well approximated by a sum of two polaron peaks. We also show in Fig. 4 the energy of the system as well as the effective interaction induced by the medium. We observe that the effective interaction follows a power law at long range and levels off at distances $\lesssim R^{\star}$. The magnitude of the effective potential of the order of $E^{\star}$ is in qualitative agreement with the numerical results Astrakharchik et al. (2023) for the weakly interacting polaron regime. Note that within the ladder approximation, the induced interaction at large distance is given by Ding et al. (2022)
@@
Conclusions. In conclusion, we have studied the ground state properties of charged polarons within the mean field approximation which in the weak coupling limit corresponding to the regularization parameter $b\gtrsim R^{\star}$ agree qualitatively with more elaborate numerical treatment. For two pinned ions, the method also provides reasonable cloud density profiles and induced interactions. This makes it promising for several future extensions. For instance, switching from static to the dynamical scenario Ardila (2021), as well as including external traps and nonzero total momentum will enable the study of the impurity effective mass and other transport properties, while Monte Carlo-based methods require a lot more intense numerical effort to probe nonequilibrium situations. Finally, it would be valuable to enrich the current ansatz to include bosonic correlations explicitly by means of Jastrow wave function Drescher et al. (2020), which may extend the validity of the method to larger coupling strength and resonant interactions within the bath.

### v2

@@ Abstract
Ionic Bose polarons are quantum entities emerging from the interaction between an ion and a Bose-Einstein condensate (BEC), featuring long-ranged interactions that can compete with the gas healing length, resulting in strong interparticle correlations and enhancement of the gas density around the ion. Full numerical treatment of such systems is computationally very expensive and does not easily allow to study the system dynamics. For this purpose, we study a mean-field based description of such systems in the co-moving frame. We find that it captures a sizeable change in the gas density and qualitatively reproduces the available results based on Monte Carlo simulations. We consider a couple of scenarios which consist of a single and two pinned ions, where it is possible to extract their effective interaction induced by the bath. This approach seems to be promising for studying transport and nonequilibrium dynamics of charged (bi)polarons in condensed media.
@@
Figure 1: The bare (solid lines) interaction given by eq. (10) and the effective potential (dotted lines) resulting from eq. (6) for several atom-ion scattering length values with $c=0.0023R^{\star}$, $n_{0}^{-1/3}=R^{\star}$ and $a_{\mathrm{B}}=0.0269R^{\star}$. The inset shows the weakly repulsive character of the effective interaction close to $R^{\star}$. The cartoon illustrates how the buildup of the density of repulsive bosons leads to the dressing of the interaction, with the upper purple (grey) line depicting the repulsive contribution to the effective potential.
@@
$\begin{split}H_{\mathrm{LLP}}&=\frac{1}{4m_{\mathrm{I}}}\left(\bm{p}_{0}-\sum_{i}\hat{\bm{p}}_{\mathrm{B},i}\right)^{2}+\frac{\hat{P}^{2}}{m_{\mathrm{I}}}+\sum_{i}\sum_{\alpha=\pm}V_{\mathrm{IB}}(\hat{\bm{x}}_{\mathrm{B},i\alpha})\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\sum_{i}\frac{\hat{\bm{p}}_{\mathrm{B},i}^{2}}{2m_{\mathrm{B}}}+\frac{1}{2}\sum_{i\neq j}V_{\mathrm{BB}}(\hat{\bm{x}}_{\mathrm{B},i}-\hat{\bm{x}}_{\mathrm{B},j}),\end{split}$ (7)
@@
Figure 2: Left: energy of a single polaron in units of $E^{\star}=1/[m_{\mathrm{B}}(R^{\star})^{2}]$ as a function of the regularization parameter $b$, with $n_{0}=0.1458\,(R^{\star})^{-3}$. Results of the modified mean field theory in the co-moving frame (straight red (grey) line) are compared with QMC calculation (green (gray) dots) as well as standard mean-field theory where $E_{1}=-\pi^{2}n/b$ (light blue (light grey) dashed) and the GPE calculation (yellow (ligt gray) dots). Right: results for higher density $n_{0}=(R^{\star})^{-3}$ with the T-matrix results of Christensen et al. (2021) (green (gray) dashed with dots) and the mean-field prediction (light blue (light grey) dashed) for comparison.
@@
The minimization of the MMF functional (4) yields the ground state energy of the system which we show in Fig. 2 and compare to the ladder approximation result of Christensen et al. (2021) using exactly the same parameters. The two approaches agree well in the limit of very weak interactions $b\gtrsim 1$ as expected. For deeper potentials, the MMF curve bends towards lower energies due to the buildup of atomic density around the ion which increases the integrated interaction energy. The standard mean field formula $E_{1}=4\pi na/m$ Scazza et al. (2022) with $a$ being the ion-atom scattering length lies very closely to the ladder approximation result as it also assumes a homogeneous gas. We also include a comparison with a QMC calculation for 192 bosons at a lower density $n_{0}=0.1458\,(R^{\star})^{-3}$. The MMF and QMC agree very well in the whole range of $b$ considered, while standard mean field theory leads to a higher estimation of energy. Also the solution of standard GPE equation for two coupled fields representing the ion and the atoms, shown as yellow dots, can only reproduce the QMC and MMF results for large values of $b$ where essentially all approaches agree with each other.
Figure 3: Upper: the gas density as a function of the distance from the impurity $r$ for several values of $b$. The dots indicate the results of solving the functional in full three-dimensional space. Lower: the number of bosons in the polaronic cloud as a function of the potential regularization parameter $b$ for one ion (light blue (grey) with dots) and two pinned ions at two different distances $R$ (straight and dashed red (grey) lines). The other parameters are the same as in Fig. 1.
@@
Figure 4: Left: Cut of the gas density profile along $z$ for $r=0$ for two pinned ions. Here $\xi_{2}=R^{\star}$, $c=0.0023R^{\star}$, $b=0.572R^{\star}$, and $n_{0}=1.9531(R^{\star})^{-3}$. The red dots indicate the results of a full three-dimensional calculation. Middle: absolute value of the system energy in units of $E^{\star}$ for two ions as a function of their separation for the same parameters and several $b$ values, right: the resulting induced effective interaction potentials. Dashed line indicates the power law derived within the ladder approximation  Ding et al. (2022) described by eq. (12).
In the case of two impurities we set up a 2D rectangular mesh, composed of 100 elements in each dimension, and second-order Lagrange polynomials for space discretization as well. Note that if we assumed $\xi_{1}=R^{\star}$, the corresponding value for the healing length is now $\xi_{2}=(\sqrt{3}/2)R^{\star}$ for the same condensate density $n_{0}$. The impurity-impurity-boson correlation density profiles, depicted in Fig. 4, reach a maximum value where the impurities are located. We further observe that the number of bound bosons is roughly twice the single-impurity case for both potentials, as presented by the red curve in the lower panel of Fig. 3. This would be expected if the induced interactions are relatively weak such that the system can be regarded as two almost independent polarons. Indeed, inspecting the density profile shows that the gas density is quite well approximated by a sum of two polaron peaks. We also show in Fig. 4 the energy of the system as well as the effective interaction induced by the medium. We observe that the effective interaction follows a power law at long range and levels off at distances $\lesssim R^{\star}$, where the buildup of the density starts being significant. The magnitude of the effective potential of the order of $E^{\star}$ is in qualitative agreement with the numerical results Astrakharchik et al. (2023) for the weakly interacting polaron regime. Note that within the ladder approximation, the induced interaction at large distance is given by Ding et al. (2022)
@@
Conclusions. In conclusion, we have studied the ground state properties of charged polarons within the mean field approximation which in the weak coupling limit corresponding to the regularization parameter $b\gtrsim R^{\star}$ agree qualitatively with more elaborate numerical treatment. For two pinned ions, the method also provides reasonable cloud density profiles and induced interactions. This makes it promising for several future extensions. For instance, switching from static to the dynamical scenario Ardila (2021); Seetharam et al. (2021), as well as including external traps and nonzero total momentum will enable the study of the impurity effective mass and other transport properties, while Monte Carlo-based methods require a lot more intense numerical effort to probe nonequilibrium situations. Note that the considered limit of long range but weak potential cannot describe the creation of many-body bound states, making the system resemble the case of neutral impurity, but with notable quantitative differences in the cloud size and density. It would thus be interesting to enrich the current ansatz to include bosonic correlations explicitly by means of Jastrow wave function Drescher et al. (2020), which may extend the validity of the method to larger coupling strength and resonant interactions within the bath.